Item 14 0.01 kg of a radioactive ligand probe with a half life of 6 hours...

Item 14

0.01 kg of a radioactive ligand probe with a half life of 6 hours and atomic mass of 96 g/mol is injected into a 90 kg patient. Each decay delivers a 3 meV α particla.

Part A - What is the equivalent dose received by the patient after all of the probe has decayed?

Dose equivalent = 350 mSv

Dose equivalent = 2.59 Sv

Dose equivalent = 6.69 Sv

Dose equivalent = 1.75 Sv

Dose equivalent = 130 mSv

Dose equivalent = 648 mSv
Answer #1

number of moles of radioactive ligand probe = 10/96 = 0.10416

number of molecules present in 0.10416 moles = 0.10416*6.023*10^23

= 6.273*10^22

since each deacy delivers a 3meV alpha particle

So total energy delivered by all the decays = 3meV * 6.273*10^22

= 1.88218*10^29 eV

= 3.015*10^10 joules

1 Sv = 1 joule/kilogram

Dose equivalent = 3.015*10^10 joules / 90 kg

= 350 mSv